Inexact Newton solvers in incremental elasto-plasticity

Radim Blaheta, Institute of Geonics

Abstract

Plasticity problems are mostly solved by incremental finite element algorithms which require solution of large nonlinear systems within each of the load steps. In dependence on the kind of the adopted stress computation procedure (explicit tangential or return mapping), these nonlinear systems can be solved by inexact Newton-like or inexact Newton methods. The main aims of our paper will be to show some recent convergence results concerning these nonlinear solvers with a discussion of their efficient implementation for the solution of 3D problems of plasticity.